# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 2.1 Ciminion
$\begin{split}\mathcal{R}^{(l)}:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},\end{split}$ (9)
where $c_{1}^{(l)},c_{2}^{(l)},c_{3}^{(l)}\in\mathbb{F}_{q}$ and $c_{4}^{(l)}\in\mathbb{F}_{q}\setminus\{0,1\}$. Note that the Feistel network $(x,y,z)^{\intercal}\mapsto(x,y,z+x\cdot y)^{\intercal}$ is also known as Toffoli gate which plays a special role in reversible computing [Tof80].
Figure 1: Encryption with Ciminion over a finite field $\mathbb{F}_{q}$.
The permutation $p_{C}$ is defined as the $r_{C}$-fold composition of $\mathcal{R}^{(l)}$, and the function $p_{E}$ is defined as the $r_{E}$-fold composition of $\mathcal{R}^{(l)}$. Finally, the rolling function is defined as
$\begin{split}rol:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}z+x\cdot y\\ x\\ y\end{pmatrix}.\end{split}$ (10)
Formulae to compute the round numbers for Ciminion are given in Table 2.
Table 2: Proposed round numbers for Ciminion [DGGK21, Table 1]. The security level $s$ must satisfy $64\leq s\leq\log_{2}\left(q\right)$, and $q\geq 2^{64}$.
Instance $p_{C}$ $p_{E}$
Standard $s+6$ $\max\left\{\left\lceil\frac{s+37}{12}\right\rceil,6\right\}$
Data limit $2^{\frac{s}{2}}$ elements $\frac{2\cdot(s+6)}{3}$ $\max\left\{\left\lceil\frac{s+37}{12}\right\rceil,6\right\}$
Conservative $s+6$ $\max\left\{\left\lceil\frac{3}{2}\cdot\frac{s+37}{12}\right\rceil,9\right\}$
E.g., for $s=128$ and data limit $2^{\frac{s}{2}}$ we have the round numbers $r_{C}=90$ and $r_{E}=14$.
It is now straight-forward to set up the iterated polynomial system for the first key pair $(K_{1},K_{2})$ given the first plain/ciphertext pair. For ease of writing we denote the $i$th intermediate state variables as $\mathbf{x}_{j}^{(i)}=\left(x_{j,1}^{(i)},x_{j,2}^{(i)},x_{j,3}^{(i)}\right)^{\intercal}$ where $j\in\{C,E\}$.
@@
Let $\mathbb{F}_{q}$ be a finite field, and let $\big{(}\aleph,(p_{1},p_{2}),(c_{1},c_{2})\big{)}\in\mathbb{F}_{q}\times\mathbb{F}_{q}^{2}\times\mathbb{F}_{q}^{2}$ be a nonce/plain/ciphertext sample given by a Ciminion encryption function. We call the polynomial system $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(i)}\right\}_{1\leq i\leq e_{C}+r_{E}}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}$ the iterated Ciminion polynomial system
$\mathbf{f}^{(i)}=\begin{dcases}\mathcal{R}^{(1)}\left(\aleph,y_{1},y_{2}\right)-\mathbf{x}_{C}^{(1)},&i=1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}_{C}^{(i-1)}\right)-\mathbf{x}_{C}^{(i)},&2\leq i\leq r_{C}-1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}_{C}^{(r_{C}-1)}\right)-\mathbf{x}_{E}^{(1)},&i=r_{C},\\ \mathcal{R}^{(i-r_{C})}\left(\mathbf{x}_{E}^{(i-r_{C})}\right)-\mathbf{x}_{E}^{(i-r_{C}+1)},&r_{C}+1\leq i\leq r_{C}+r_{E}-1,\\ \mathcal{R}^{(r_{E})}\left(\mathbf{x}_{E}^{(r_{E})}\right)-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix},&i=r_{C}+r_{E}.\end{dcases}$
@@ 2.2 Hydra
Let $\mathbb{F}_{q}$ be a finite field, the Hydra [GØSW23] PRF over $\mathbb{F}_{q}$ is most also easily described via an illustration, see Figure 2. A nonce $\aleph\in\mathbb{F}_{q}$ and an initial value $\texttt{IV}\in\mathbb{F}_{q}^{3}$ are added to a key $\texttt{K}\in\mathbb{F}_{q}^{4}$, and then fed into the keyed body function $\mathcal{B}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$ which produces the state $\mathbf{y}||\mathbf{z}=(\mathbf{y},\mathbf{z})^{\intercal}\in\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}$. Then $\mathbf{y}||\mathbf{z}$ is fed into the first keyed head function $\mathcal{H}_{\texttt{K}}:\mathbb{F}_{q}^{8}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$. After adding $\mathbf{y}||\mathbf{z}$ to the output of the head the result is released as the first Hydra sample. Say one has already requested $(i-1)$ Hydra samples, for an additional sample $\mathbf{y}||\mathbf{z}$ is fed into the $i$th rolling function $\mathcal{R}_{i}:\mathbb{F}_{q}^{8}\to\mathbb{F}_{q}^{8}$, and then again into the head $\mathcal{H}_{\texttt{K}}$. After adding the output of $\mathcal{R}_{i}$ to the output of $\mathcal{H}_{\texttt{K}}$ the result is released as $i$th Hydra sample.
@@
For our polynomial model we are only going to consider the heads of the Hydra. Thus, we do not describe the details of the body function $\mathcal{B}$, interested readers can find the description in [GØSW23, §5.2]. Let $\mathbf{M}_{\mathcal{E}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be an MDS matrix, e.g. the Hydra designers recommend $\operatorname{circ}\left(2,3,1,1\right)$ or $\operatorname{circ}\left(3,2,1,1\right)$222 We consider circulant matrices with respect to a right shift, i.e.  $\operatorname{circ}(a_{1},\dots,a_{n})=\begin{pmatrix}a_{1}&a_{2}&\dots&a_{n-1}&a_{n}\\ a_{n}&a_{1}&\dots&a_{n-2}&a_{n-1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ a_{2}&a_{3}&\dots&a_{n}&a_{1}\end{pmatrix}.$ For the keyed permutation $\mathcal{H}_{\texttt{K}}$, let
@@
$f(\mathbf{x})$ $=\left(\sum_{i=1}^{8}(-1)^{\left\lfloor\frac{i-1}{4}\right\rfloor}\cdot x_{i}\right)^{2}.$ (12)
Moreover, let $\mathbf{M}_{\mathcal{J}}\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$ be an invertible matrix. Then the heads round function is defined as
@@
Let $\mathbf{M}_{\mathcal{I}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be a matrix of the form
@@
Finally, for the rolling function let
@@
$\begin{split}\mathcal{R}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}&\mapsto\mathbf{M}_{\mathcal{R}}\begin{pmatrix}\{y_{i}+g(\mathbf{y},\mathbf{z})\}_{1\leq i\leq 4}\\ \{z_{i}+g(\mathbf{z},\mathbf{y})\}_{1\leq i\leq 4}\end{pmatrix},\end{split}$ (18)
and the $i$th rolling function is simply the $i$-fold composition $\mathcal{R}_{i}=\mathcal{R}\circ\cdots\circ\mathcal{R}$. Note that the degree increasing parts in $\mathcal{J}_{i}$ and $\mathcal{R}$ are generalized Lai–Massey permutations [RS22].
@@
Given two Hydra outputs it is straight-forward to set up an overdetermined polynomial system for Hydra. For ease of writing we denote the $i$th intermediate state variables for the $j$th output as $\mathbf{x}^{(i)}_{j}=\left(x_{j,1}^{(i)},\dots,x_{j,8}^{(i)}\right)^{\intercal}$ and analog for $\mathbf{y}=(y_{1},\dots,y_{4})^{\intercal}$, $\mathbf{z}=(z_{1},\dots,z_{4})^{\intercal}$ and $\mathbf{k}=(k_{1},\dots,k_{4})^{\intercal}$.
@@
$\mathbf{f}_{\mathcal{R}}$ $=\mathcal{R}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}\right)-\mathbf{x}_{2}^{(0)},$
$\mathbf{f}_{2}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\mathbf{x}_{2}^{(0)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{2}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{2}^{(r_{\mathcal{H}})},\mathbf{k}^{\prime}\right)+\mathbf{x}_{2}^{(1)}-\mathbf{c}_{2},&i=r_{\mathcal{H}},\end{dcases}$
@@
This polynomial system consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables. It can also easily be expanded to an arbitrary number of output samples.
@@ 2.3 Gröbner Bases
Now let $I\subset P=K[x_{1},\dots,x_{n}]$ be an ideal, and let $f\in P$. By $\operatorname{LM}_{>}\left(f\right)$ we denote the leading monomial of $f$, i.e. the largest monomial with non-zero coefficient in $f$ under $>$. A finite ideal basis $\mathcal{G}=\{g_{1},\dots,g_{m}\}\subset I$ such that
$\big{(}\operatorname{LM}_{>}(g_{1}),\dots,\operatorname{LM}_{>}(g_{m})\big{)}=\operatorname{in}_{>}(I)$ (20)
is called a $>$-Gröbner basis of $I$. I.e., the leading monomials of $\mathcal{G}$ are minimal under $>$. With Gröbner bases one can solve many important computational problems like the membership problem or the computation of the zero locus. For a general introduction to Gröbner bases we refer to [KR00, KR05, CLO15].
@@ 2.3.1 Linear Algebra-Based Gröbner Basis Algorithms
The first algorithm to compute Gröbner bases was Buchberger’s algorithm, see e.g. [CLO15, Chapter 2 §7]. Given an input polynomial system $\mathcal{F}\subset P=k[x_{1},\dots,x_{n}]$ and a term order $>$ on $P$, Buchberger’s algorithms performs pairwise cancellation of the leading monomials under $>$ of two elements in $\mathcal{F}$ followed by multivariate division by remainder. If a non-zero element is produced, then it is added to $\mathcal{F}$ and the process is repeated. Buchberger proved that this procedure terminates and produces a $>$-Gröbner basis.
Lazard [Laz83] generalized this construction by performing multiple leading term cancellations and reductions at the same time via a matrix representation of the polynomial system $\mathcal{F}$. Assume that =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset P$ is homogeneous and fix a term order $>$ on $P$. The homogeneous Macaulay matrix $M_{d}$ for $\mathcal{F}$ has columns indexed by the monomials of degree $d$ in $P$ sorted via $>$, and its rows are indexed by $s\cdot f$, where $f\in\mathcal{F}$ and $s\in P$ is a monomial such that $\deg\left(s\cdot f\right)=d$. The entry of row $s\cdot f$ at column $t$ is the coefficient of $t$ in the polynomial $s\cdot f$. If $\mathcal{F}$ is inhomogeneous, then one replaces the degree equalities by inequalities to obtain the inhomogeneous Macaulay matrix $M_{\leq d}$.
Obviously, if we perform Gaussian elimination on $M_{0},\dots,M_{d}$ respectively $M_{\leq d}$ for a large enough value of $d$, then we construct a $>$-Gröbner basis for $\mathcal{F}$. The least such $d$ is called the solving degree, to the best of our knowledge this notion was first introduced in [DS13].
@@
If $\mathcal{F}$ is homogeneous, we consider the homogeneous Macaulay matrix $M_{d}$ and let the solving degree of $\mathcal{F}$ be the least degree $d$ such that Gaussian elimination on $M_{0},\dots,M_{d}$ produces a Gröbner basis of $\mathcal{F}$ with respect to $>$.
By [CG21, Theorem 7] the homogeneous solving degree always upper bounds the inhomogeneous one, i.e.
$\operatorname{sd}_{DRL}\left(\mathcal{F}\right)\leq\operatorname{sd}_{DRL}\left(\mathcal{F}^{\text{hom}}\right).$ (21)
Algorithms which perform Gaussian elimination on the Macaulay matrices are generally referred to as linear algebra-based Gröbner basis algorithms. Modern examples of such algorithms are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms. Since the solving degree is a priori unknown for most polynomial system, F4/5 construct the Macaulay matrices and perform Gaussian elimination for increasing values of $d$. Therefore, they require a stopping criterion to decide whether a Gröbner basis has already been found. On the other hand, we can always stop F4/5 after the final Gaussian elimination in the solving degree, since the row space basis must already contain a Gröbner basis. For this reason, we also consider the solving degree as complexity measure for advanced linear algebra-based algorithms.
2.3.2 Complexity Estimates via the Solving Degree
It is well-known that the complexity of Gaussian elimination of a matrix $\mathbf{A}\in K^{n\times m}$, where $K$ is a field and $r=\operatorname{rank}\left(\mathbf{A}\right)$, is given by [Sto00, §2.2] $\mathcal{O}\left(n\cdot m\cdot r^{\omega-2}\right)$ field operations, where $2.372862\leq\omega<$ 2.37286 is a linear algebra constant [AW21].
Given a homogeneous polynomial system =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset P=K[x_{1},\dots,x_{n}]$, it is well-known that the number of monomials in $P$ of degree $d$ is $\binom{n+d-1}{d}$. Moreover, for $f\in\mathcal{F}$ the number of rows $\hat{m}\cdot f$, where $\hat{m}\in P$ a monomial such that $\deg\left(\hat{m}\cdot f\right)=d$, is also bounded by $\binom{n+d-1}{d}$. Thus, the complexity of Gaussian elimination on the homogeneous Macaulay matrices $M_{0},\dots,M_{d}$ is bounded by
$\mathcal{O}\left(m\cdot d\cdot\binom{n+d-1}{d}^{\omega}\right).$ (22)
Also, note that dehomogenization of a Gröbner basis yields an inhomogeneous Gröbner basis [KR05, Proposition 4.3.18] (if the term order is compatible with homogenization). Therefore, we can also regard Equation 22 as complexity estimate for Gaussian elimination on the inhomogeneous Macaulay matrix $M_{\leq d}$.
2.4 Generic Coordinates
Caminata & Gorla [CG23] have proven that for a special class of polynomial systems the DRL solving degree is always upper bounded by the so-called Macaulay bound. They called these polynomial to be in generic coordinates. For the formal definition of this notion, we use the shorthand notation “$r\nmid 0\mod I$” for expressing that $r$ is not a zero-divisor in $P/I$. Moreover, we always assume that the homogenization variable $x_{0}$ is the least variable under DRL.
Definition 2.4 ([CG21, CG22, Definition 5]).
Let $K$ be an infinite field. Let $I\subset K[x_{0},\dots,\allowbreak x_{n}]$ be a homogeneous ideal with $|\mathcal{Z}_{+}(I)|<\infty$. We say that $I$ is in generic coordinates if either $|\mathcal{Z}_{+}(I)|=0$ or $x_{0}\nmid 0\mod I^{\text{sat}}$.

### v4

@@ 2.1 Ciminion
$\begin{split}\mathcal{R}^{(i)}:\mathbb{F}_{q}^{3}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y\end{pmatrix}+\begin{pmatrix}c_{1}^{(i)}\\ c_{2}^{(i)}\\ c_{3}^{(i)}\end{pmatrix},\end{split}$ (2)
where $c_{1}^{(i)},c_{2}^{(i)},c_{3}^{(i)}\in\mathbb{F}_{q}$ and $c_{4}^{(i)}\in\mathbb{F}_{q}\setminus\{0,1\}$. Note that the Feistel network $(x,y,z)^{\intercal}\mapsto(x,y,z+x\cdot y)^{\intercal}$ is also known as Toffoli gate which plays a special role in reversible computing [Tof80]. Finally, the rolling function is defined as
$rol:\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3},\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto\begin{pmatrix}z+x\cdot y\\ x\\ y\end{pmatrix}.$ (3)
Figure 1: Encryption with Ciminion.
Formulas to compute the round numbers for Ciminion are given in [DGGK21, Table 1]. E.g., for $s=128$ and data limit $2^{\frac{s}{2}}$ we have the round numbers $r_{C}=90$ and $r_{E}=14$.
It is now straight-forward to set up the iterated polynomial system for the first key pair $(K_{1},K_{2})$ given the first plain/ciphertext pair. For ease of notation we abbreviate $r=r_{C}+r_{E}$.
@@
Let $\mathbb{F}_{q}$ be a finite field, and let $\big{(}\aleph,(p_{1},p_{2}),(c_{1},c_{2})\big{)}\in\mathbb{F}_{q}\times\mathbb{F}_{q}^{2}\times\mathbb{F}_{q}^{2}$ be a nonce/plain/ciphertext sample given by a Ciminion encryption function. We call the polynomial system $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(i)}\right\}_{1\leq i\leq r}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(r-1)},x\Big{]}$ the iterated Ciminion polynomial system
$\mathbf{f}^{(i)}=\begin{dcases}\mathcal{R}^{(1)}\left(\aleph,y_{1},y_{2}\right)-\mathbf{x}^{(1)},&i=1,\\ \mathcal{R}^{(i)}\left(\mathbf{x}^{(i-1)}\right)-\mathbf{x}^{(i)},&2\leq i\leq r-1,\\ \mathcal{R}^{(r)}\left(\mathbf{x}^{(r-1)}\right)-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix},&i=r.\end{dcases}$
@@ 2.2 Hydra
Let $\mathbb{F}_{q}$ be a finite field, the Hydra [GØSW23] PRF over $\mathbb{F}_{q}$ is also easily described via an illustration, see Figure 2. A nonce $\aleph\in\mathbb{F}_{q}$ and an initial value $\texttt{IV}\in\mathbb{F}_{q}^{3}$ are added to a key $\texttt{K}\in\mathbb{F}_{q}^{4}$, and then fed into the keyed body function $\mathcal{B}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$ which produces the state $\mathbf{y}||\mathbf{z}=(\mathbf{y},\mathbf{z})^{\intercal}\in\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}$. Then $\mathbf{y}||\mathbf{z}$ is fed into the first keyed head function $\mathcal{H}_{\texttt{K}}:\mathbb{F}_{q}^{8}\times\mathbb{F}_{q}^{4}\to\mathbb{F}_{q}^{8}$. After adding $\mathbf{y}||\mathbf{z}$ to the output of the head the result is released as the first Hydra sample. Say one has already requested $(i-1)$ Hydra samples, for an additional sample $\mathbf{y}||\mathbf{z}$ is fed into the $i$th rolling function $\mathcal{R}_{i}:\mathbb{F}_{q}^{8}\to\mathbb{F}_{q}^{8}$, and then again into the head $\mathcal{H}_{\texttt{K}}$. After adding the output of $\mathcal{R}_{i}$ to the output of $\mathcal{H}_{\texttt{K}}$ the result is released as $i$th Hydra sample.
@@
For our polynomial model we are only going to consider the heads of the Hydra. Thus, we do not describe the details of the body function $\mathcal{B}$, interested readers can find the description in [GØSW23, §5.2]. Let $\mathbf{M}_{\mathcal{E}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be an MDS matrix, e.g. the Hydra designers recommend $\operatorname{circ}\left(2,3,1,1\right)$ or $\operatorname{circ}\left(3,2,1,1\right)$.333 We consider circulant matrices with respect to a right shift, i.e.  $\operatorname{circ}(a_{1},\dots,a_{n})=\begin{pmatrix}a_{1}&a_{2}&\ldots&a_{n-1}&a_{n}\\ a_{n}&a_{1}&\ldots&a_{n-2}&a_{n-1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ a_{2}&a_{3}&\ldots&a_{n}&a_{1}\end{pmatrix}.$ For the keyed permutation $\mathcal{H}_{\texttt{K}}$, let
@@
$f(x_{1},\dots,x_{n})$ $=\left(\sum_{i=1}^{8}(-1)^{\left\lfloor\frac{i-1}{4}\right\rfloor}\cdot x_{i}\right)^{2}.$ (5)
Moreover, let $\mathbf{M}_{\mathcal{J}}\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$ be an invertible matrix. The heads round function is defined as
@@
Let $\mathbf{M}_{\mathcal{I}}\in\operatorname{GL}_{4}\left(\mathbb{F}_{q}\right)$ be an invertible matrix of the form
@@
Note that the $\mu_{i,j}\in\mathbb{F}_{q}$ have to be chosen under additional constraints to prevent invariant subspaces [GØSW23, §5.2]. Finally, for the rolling function let
@@
$\begin{split}\mathcal{R}:\mathbb{F}_{q}^{4}\times\mathbb{F}_{q}^{4}&\to\mathbb{F}_{q}^{8},\\ \begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}&\mapsto\mathbf{M}_{\mathcal{R}}\begin{pmatrix}\big{(}y_{i}+g(\mathbf{y},\mathbf{z})\big{)}_{1\leq i\leq 4}\\ \big{(}z_{i}+g(\mathbf{z},\mathbf{y})\big{)}_{1\leq i\leq 4}\end{pmatrix},\end{split}$ (11)
and the $i$th rolling function is simply the composition $\mathcal{R}_{i}=\mathcal{R}\circ\mathcal{R}_{i-1}+\mathbf{c}_{\mathcal{R}}^{(i)}$, where $\mathbf{c}_{\mathcal{R}}^{(i)}\in\mathbb{F}_{q}$ is a random constant. Note that the degree increasing parts in $\mathcal{J}_{i}$ and $\mathcal{R}$ are generalized Lai–Massey permutations [GOPS22, RS24].
@@
Given two Hydra outputs it is straight-forward to set up an overdetermined polynomial system for Hydra. For ease of writing we denote the $i$th intermediate state variables for the $j$th output as $\mathbf{x}^{(i)}_{j}=\left(x_{j,1}^{(i)},\dots,x_{j,8}^{(i)}\right)^{\intercal}$ and analogously for $\mathbf{y}=(y_{1},\dots,y_{4})^{\intercal}$, $\mathbf{z}=(z_{1},\dots,z_{4})^{\intercal}$ and $\mathbf{k}=(k_{1},\dots,k_{4})^{\intercal}$.
@@
$\mathbf{f}_{\mathcal{R}}$ $=\mathcal{R}\left(\begin{pmatrix}\mathbf{y}\\ \mathbf{z}\end{pmatrix}\right)+\mathbf{c}_{\mathcal{R}}^{(1)}-\mathbf{x}_{2}^{(0)},$
$\mathbf{f}_{2}^{(i)}$ =$=\begin{dcases}\mathcal{J}_{1}\left(\mathbf{x}_{2}^{(0)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(1)},&i=1,\\ \mathcal{J}_{i}\left(\mathbf{x}_{2}^{(i-1)},\mathbf{k}^{\prime}\right)-\mathbf{x}_{2}^{(i)},&2\leq i\leq r_{\mathcal{H}}-1,\\ \mathcal{J}_{r_{\mathcal{H}}}\left(\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}^{\prime}\right)+\mathbf{x}_{2}^{(0)}-\mathbf{c}_{2},&i=r_{\mathcal{H}},\end{dcases}$
@@
This polynomial system consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables. It can also easily be extended to an arbitrary number of output samples.
@@ 2.3 Gröbner Bases
Now let $I\subset P=K[x_{1},\dots,x_{n}]$ be an ideal, and let $>$ be a term order on $P$. For $f\in P$ we denote with $\operatorname{LM}_{>}\left(f\right)$ the leading monomial of $f$, i.e. the largest monomial with non-zero coefficient in $f$ under $>$. A finite ideal basis $\mathcal{G}=\{g_{1},\dots,g_{m}\}\subset I$ such that
$\big{(}\operatorname{LM}_{>}(g_{1}),\dots,\operatorname{LM}_{>}(g_{m})\big{)}=\big{(}\operatorname{LM}_{>}(f)\;\big{|}\;f\in I\big{)}$ (13)
is called a $>$-Gröbner basis of $I$. I.e., the leading monomials of $\mathcal{G}$ are minimal under $>$.
We abbreviate the ideal of leading terms of $I$ as $\operatorname{LM}_{>}(I)=\big{(}\operatorname{LM}_{>}(f)\;\big{|}\;f\in I\big{)}$. Also, for ease of writing we often denote the variables of $P$ in vector form $\mathbf{x}=(x_{1},\dots,x_{n})^{\intercal}$. If not specified otherwise, then we always assume that the variables within a variable vector $\mathbf{x}$ are naturally ordered with respect to $>$, i.e. $x_{1}>\ldots>x_{n}$.
With Gröbner bases one can solve many important computational problems like the membership problem or the computation of the set of zeros of an ideal. For a general introduction to Gröbner bases we refer to [KR00, KR05, CLO15].
In this paper we work with the lexicographic (LEX) and the degree reverse lexicographic (DRL) term order. For $\mathbf{a},\mathbf{b}\in\mathbb{Z}_{\geq 0}^{n}$, we have that $\mathbf{a}>_{LEX}\mathbf{b}$ if the first non-zero entry of $\mathbf{a}-\mathbf{b}$ is positive. We have that $\mathbf{a}>_{DRL}\mathbf{b}$ if either $\sum_{i=1}^{n}a_{i}>\sum_{i=1}^{n}b_{i}$ or $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}b_{i}$ and the last non-zero entry of $\mathbf{a}-\mathbf{b}$ is negative.
In more generality, a term order $>$ on $P$ is called degree compatible if for any $f,g\in P$ the degree inequality $\deg\left(f\right)>\deg\left(g\right)$ implies that $f>g$. Obviously, the DRL term order is degree compatible.
@@ 2.3.1 Linear Algebra-Based Gröbner Basis Algorithms
Gröbner bases can be constructed via Gaussian elimination on the so-called Macaulay matrices. Let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be homogeneous, and let $>$ be a term order on $P$. The homogeneous Macaulay matrix $M_{d}$ for $\mathcal{F}$ has columns indexed by the monomials of degree $d$ in $P$ sorted via $>$, and its rows are indexed by $s\cdot f$, where $f\in\mathcal{F}$ and $s\in P$ is a monomial such that $\deg\left(s\cdot f\right)=d$. The entry of row $s\cdot f$ at column $t$ is the coefficient of $t$ in the polynomial $s\cdot f$. If $\mathcal{F}$ is inhomogeneous, then one replaces the degree equalities by inequalities to obtain the inhomogeneous Macaulay matrix $M_{\leq d}$. Obviously, if we perform Gaussian elimination on $M_{0},\dots,M_{d}$ (respectively $M_{\leq d}$) for $d$ large enough, then we construct a $>$-Gröbner basis for $\mathcal{F}$. This idea can be traced back to Lazard [Laz83]. The least $d$ which yields a Gröbner basis is of special interest.
@@
If $\mathcal{F}$ is homogeneous, then we consider the homogeneous Macaulay matrix $M_{d}$ and let the solving degree of $\mathcal{F}$ be the least degree $d$ such that Gaussian elimination on $M_{0},\dots,M_{d}$ produces a Gröbner basis of $\mathcal{F}$ with respect to $>$.
Algorithms which perform Gaussian elimination on the Macaulay matrices are generally referred to as linear algebra-based Gröbner basis algorithms. Modern examples of such algorithms are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms. Since the solving degree is a priori unknown for most polynomial systems, F4/5 construct the Macaulay matrices for increasing values of $d$. Therefore, they require a stopping criterion to decide whether a Gröbner basis has already been found. On the other hand, we can always stop F4/5 after the final Gaussian elimination in the solving degree, since the row space basis must already contain a Gröbner basis. For this reason, we also consider the solving degree as complexity measure for advanced linear algebra-based algorithms.
2.3.2 Polynomial Systems in Generic Coordinates
Polynomial systems in generic coordinates have a proven DRL solving degree upper bound. For their definition we recall the degree of regularity of a polynomial system.
Definition 2.4 (Degree of regularity, [BFS04, Definition 4]).
Let $K$ be a field, and let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$. If $\dim\left(\mathcal{F}^{\text{top}}\right)=0$, then the degree of regularity is defined as
$d_{\operatorname{reg}}\left(\mathcal{F}\right)=\min\left\{d\geq 0\;\middle|\;\dim_{\mathbb{F}_{q}}\Big{(}\big{\{}f\in\left(\mathcal{F}^{\text{top}}\right)\;\big{|}\;\deg\left(f\right)=d\big{\}}\Big{)}=\binom{n+d-1}{d}\right\}.$
If $\dim\left(\mathcal{F}^{\text{top}}\right)>0$, then $d_{\operatorname{reg}}\left(\mathcal{F}\right)=\infty$.
We present an equivalent definition for generic coordinates which requires slightly less knowledge about commutative algebra and projective geometry than the original definition from [CG21, CG22, Definition 5].
Definition 2.5 (cf. [CG21, CG22, Definition 5], [Ste24, Theorem 3.2]).
Let $K$ be an infinite field, and let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$. We say that $\left(\mathcal{F}^{\text{hom}}\right)\subset P[x_{0}]$ is in generic coordinates if either $d_{\operatorname{reg}}\left(\mathcal{F}\right)<\infty$ or $\sqrt{\mathcal{F}^{\text{hom}}}=(x_{0},\dots,x_{n})$.